# Explainable Metrics for the Assessment of Neurodegenerative Diseases through Handwriting Analysis

### v2

@@ III-E1 Grouped metrics
III-E2 Difference between metrics
To study the variations between tasks and how changes of behavior might be signs of NDs, we compared the metrics extracted between 2 tasks. For each session, if a given participant had successfully completed both tasks of a pair, we took the difference between 2 metrics, and perform the same statistical analysis as described in section III-D for this new ’coupled’ metric. The considered couples are the following:
  • Dominant vs. Non-Dominant: The difference when drawing with their non-dominant hand is computed to measure eventual dissymmetry in their response. This difference is measured for the following tasks:
    • Point Dominant vs Non Dominant
    • Spiral Dominant vs Non Dominant
  • Reading vs Non-Reading: The difference when performing multiple tasks at the same time is measured by comparing the following tasks:
    • Point Dominant vs Point Sustained
    • Spiral Dominant vs Spiral PaTaKa
    • Copy Text vs Copy Read Text
  • Memory vs Non-Memory: The difference when a task from memory or from a given image is measured by comparing the task Copy Image against the task Copy Image Memory.
This gave us 6 new potential comparisons.
III-F Classification Experiments
In addition to conducting statistical analyses, we conduct binary classification experiments across various experimental groups using metrics extracted from each task. Namely, it involves a binary classification (e.g., AD or CTL) that facilitates the assignment of samples to their corresponding groups, thereby enabling the evaluation of the metrics’ effectiveness in distinguishing between the categories under investigation.
These experiments further evaluate the discriminative power of the metrics and the utility of tasks in extracting valuable information for characterizing different disorders. We are employing a range of classifiers, including Support Vector Machine (SVM), K-Nearest Neighbors (KNN), Random Forest (RF), Extreme Gradient Boosting (XGBoost), Bagging (BG), and Multi-Layer Perceptron (MLP)  [^4] To implement the different classifiers, we are using the Scikit-learn (Sklearn) library in Python. See https://www.tutorialspoint.com/scikit_learn/scikit_learn_introduction.htm.
Nested cross-validation (NCV) is a widely used approach for assessing model performance, especially when dealing with limited data, as it helps provide an unbiased estimate of how well the model will generalize to new data. By incorporating two levels of data partitioning, NCV enables both model evaluation and hyperparameter tuning while minimizing overfitting and overly optimistic performance estimates [42, 43]. NCV is structured around two sequential processes:
  1. 1.
    Outer loop: The data is divided into folds, where one fold serves as a test set, and the remaining folds form the training set. This is repeated so that each fold acts as the test set once, yielding a balanced estimate of model accuracy.
  2. 2.
    Inner loop: Within each outer training set, the data undergoes another round of splitting to optimize hyperparameters and select features. This tuning is isolated from the test set, ensuring the performance evaluation remains unbiased.
This separation of tuning and testing phases helps NCV deliver a realistic and robust measure of the model’s performance, making it especially effective for tasks where data scarcity could lead to overfitting.
@@ IV-A1 AD vs CTL
Over the 54 task-agnostic metrics, 48 are significant ($p\leq 0.05$) in comparing AD and CTL participants for at least one task, and 17 have p-values smaller than 0.001. We present the 10 metrics having the lowest p-values in Table III across all tasks. Our most significant and consistent metrics across all tasks are the Shannon entropy, the duration, the mean of the speed and accelerations for point tasks, and the variations of angular speed and acceleration for the drawing and writing tasks. We show the box plots for some of those metrics in Figure 10.
Shannon entropy is a measure of the uncertainty in a signal, showing the erratic aspect of one’s handwriting, which is a marker of indecision for cognitively impaired participants. The differences measured in linear speed and linear acceleration, as well as duration, all show how significantly slower AD participants are from the control group, presumably due to processing inefficiency. Finally, the differences in angular speed and angular acceleration are indicators of the swiftness of the writing. For detecting AD, the most significant metrics were all for writing and drawing tasks, which are much more demanding cognitively than the rest of tasks and mainly consisting of tasks designed for AD assessment, as opposed to the point and spiral tasks.
@@
A limitation of our approach here is the lower number of participant for the AD/MCI category, which did not allowed us a separated analysis of the cognitive impairments against the participants presenting only AD’s signs.
@@
in-air avg. angular acc. 0.006 0.700 $<$0.001 0.791 0.010 0.672
in-air shannon entropy $<$0.001 0.873 $<$0.001 0.887 $<$0.001 0.827
in-air std angular acc. $<$0.001 0.737 $<$0.001 0.818 0.003 0.702
in-air std angular speed $<$0.001 0.730 $<$0.001 0.826 0.005 0.692
on-tablet duration $<$0.001 0.801 $<$0.001 0.738 $<$0.001 0.751
on-tablet avg. speed 0.007 0.693 0.185 0.614 $<$0.001 0.784
@@
in-air avg. angular acc. 0.025 0.691 $<$0.001 0.761 $<$0.001 0.760
in-air shannon entropy $<$0.001 0.794 $<$0.001 0.866 $<$0.001 0.769
in-air std angular acc. 0.014 0.727 $<$0.001 0.768 $<$0.001 0.821
in-air std angular speed 0.025 0.694 $<$0.001 0.799 $<$0.001 0.774
on-tablet duration 0.036 0.681 $<$0.001 0.740 0.148 0.599
on-tablet avg. speed 0.346 0.602 $<$0.001 0.787 0.003 0.698
AD vs CTL Point Dom. Point Non Dom. Point Sustained
@@
in-air avg. angular acc. $<$0.001 0.726 0.077 0.627 $<$0.001 0.740
in-air avg. acc. $<$0.001 0.792 $<$0.001 0.850 $<$0.001 0.809
in-air avg. speed $<$0.001 0.797 $<$0.001 0.857 $<$0.001 0.827
in-air shannon entropy $<$0.001 0.865 $<$0.001 0.810 $<$0.001 0.872
in-air std angular acc. $<$0.001 0.804 0.003 0.708 $<$0.001 0.789
in-air std angular speed 0.021 0.665 0.131 0.608 0.112 0.615
@@ IV-A2 PD vs CTL
Over the 54 task-agnostic metrics, 19 are significant ($p\leq 0.05$) in the comparison between PD and CTL participants for at least one task, and 4 have p-values under 0.001. Table IV shows the top 10 metrics with the lowest p-values across all tasks for the PD vs. CTL comparison.
@@
PD vs CTL All Drawings Copy Image Copy Image Memory
@@
in-air avg. angular acc. 0.038 0.634 0.289 0.636 0.301 0.621
in-air avg. acc. 0.038 0.632 0.289 0.647 0.301 0.622
in-air avg. speed 0.038 0.622 0.289 0.635 0.359 0.610
in-air std angular acc. 0.009 0.670 0.227 0.679 0.301 0.659
in-air std angular speed 0.038 0.628 0.321 0.611 0.301 0.633
PD vs CTL Point Dominant Point Non-Dominant All Points
p-val AUC p-val AUC p-val AUC
in-air avg. angular acc. 0.002 0.723 $<$0.001 0.768 $<$0.001 0.723
in-air avg. acc. 0.189 0.611 0.116 0.624 0.007 0.611
in-air avg. hor. acc. 0.157 0.628 0.011 0.694 $<$0.001 0.637
in-air avg. vert. acc. 0.022 0.676 0.067 0.647 $<$0.001 0.650
in-air avg. angular speed $<$0.001 0.743 $<$0.001 0.768 $<$0.001 0.742
in-air avg. speed 0.169 0.620 0.106 0.630 0.008 0.608
in-air avg. hor. speed 0.169 0.616 0.019 0.678 $<$0.001 0.639
in-air avg. vert. speed 0.004 0.711 0.019 0.678 $<$0.001 0.682
in-air std angular acc. $<$0.001 0.738 $<$0.001 0.770 $<$0.001 0.723
in-air std angular speed $<$0.001 0.756 $<$0.001 0.759 $<$0.001 0.733
These metrics primarily revolve around in-air dynamics, variations, and averages of speeds and angular speeds. While notably discriminatory for point tasks, they also exhibit relevance, albeit to a lesser extent, for writing tasks like Copy Read Text and Freewrite. Additionally, we present box plots for these significant metrics in Figure 11 for further visualization and analysis.
@@
For detecting PD, the most significant metrics were for drawing and point tasks. The point tasks were better suited to assess motoric features of PD, such as tremor and bradykinesia. Accordingly, we observed subjects with PD showed significantly lower angular speed and angular acceleration in air, which could be indicators of hypometria [44]. The drawing tasks were much more complex tasks, mainly designed for cognitive evaluation but that still require basic motor skills to complete. In those tasks, the in-air angular speed is still a discriminative metric, but metrics of variation (as standard deviation or amplitudes) of the speed and acceleration on the tablet were notable. These findings may represent, tremor, dystonia, dyskinesia, myoclonus other other complicating motor symptoms that develop in some persons with PD [11]. However, none of the task-agnostic metrics measured on the spirals tasks, mainly motor tasks often used for PD assessment, were discriminative enough.
@@ IV-A3 PD vs PDM
Among the 54 task-agnostic metrics, 30 are significant ($p\leq 0.05$) in comparing PD and PDM participants for at least one task. We present the 10 metrics having the lowest p-values in Table V across all tasks. We only find significant metrics for Point and Spiral tasks here.
To distinguish PD from other parkinsonism, significant variations can be found in the Point and Spiral tasks and motor exercises designed for PD assessment. We observe few significant metrics for the Point tasks, mostly amplitudes of speed and acceleration, the best being for vertical velocity. PD participants have lower variations in their speed than PDM participants, which are still lower than healthy controls, as we can see in Figure 12. The Spiral tasks reveal a difference in speed between PD and PDM participants, the latest being slower, as shown in Figure 13.
TABLE V: $p$-values and AUC for the best 10 metrics for PD vs PDM comparison. The table is repeated to fit the page. Significant p-values are bolded, as well as AUC values over 0.75.
PD vs PDM Spiral Dominant Spiral Pataka All Spirals
p-val AUC p-val AUC p-val AUC
on-tablet duration 0.323 0.680 0.128 0.686 0.003 0.685
on-tablet avg. angular acc. 0.323 0.696 0.128 0.726 0.003 0.687
on-tablet avg. hor. acc. 0.323 0.662 0.128 0.721 0.005 0.676
on-tablet avg. vert. acc. 0.323 0.662 0.128 0.710 0.003 0.685
on-tablet avg. hor. speed 0.323 0.676 0.128 0.736 0.003 0.696
on-tablet avg. vert. speed 0.323 0.700 0.128 0.752 $<$0.001 0.724
on-tablet shannon entropy 0.323 0.662 0.128 0.698 0.006 0.671
on-tablet std vert. speed 0.398 0.649 0.128 0.719 0.003 0.685
PD vs PDM Point Dominant Point Sustained All Points
p-val AUC p-val AUC p-val AUC
in-air amp vert. acc. 0.475 0.626 0.042 0.800 0.004 0.701
in-air amp vert. speed 0.540 0.607 0.063 0.767 0.005 0.690
Figure 12: Boxplots of some the best metrics extracted for PD vs PDM on the point tasks.
@@ IV-B1 Point Tasks Results
Figure 14 shows all metrics with their significance. None of the metrics are consistently significant across all tasks, but the radius average and radius standard deviation metrics exhibit the same behavior across all tasks: metrics from AD participants are higher than in the other groups. We observe a slightly higher radius average for PD participants compared to CTL, only significant when considering all point tasks, showing that the Point tasks capture some of the expected tremors.
@@ IV-B2 Spiral Tasks Results
TABLE VI: $p$-values and AUC of the specific metrics for all spiral tasks significant for at least one comparison. Significant p-values and AUC over 0.75 are bolded.
Spiral Metrics AD vs CTL PD vs CTL PD vs PDM
@@
N loops 0.092 0.573 $<$0.001 0.604 0.303 0.538
All avg radius diff 0.141 0.557 0.019 0.568 0.945 0.554
Spirals Duration $<$0.001 0.695 0.104 0.543 0.003 0.623
N Loops 0.370 0.563 $<$0.001 0.615 0.209 0.542
Fund. Period $<$0.001 0.694 0.109 0.545 0.004 0.615
Spiral Duration $<$0.001 0.695 0.496 0.543 0.155 0.623
Dom. Fund. Period $<$0.001 0.694 0.474 0.545 0.185 0.615
Spiral N loops 0.714 0.573 0.018 0.604 0.166 0.538
Non Duration 0.123 0.695 0.043 0.543 0.028 0.623
Dom. N Loops 0.674 0.563 0.031 0.615 0.241 0.542
Fund. Period 0.100 0.694 0.051 0.545 0.035 0.615
N loops 0.150 0.573 0.012 0.604 0.429 0.538
Spiral Duration $<$0.001 0.695 0.700 0.543 0.105 0.623
Pataka N Loops 0.406 0.563 0.007 0.615 0.620 0.542
Fund. Period $<$0.001 0.694 0.658 0.545 0.103 0.615
@@ IV-B3 Numbers Task Results
TABLE VII: $p$-values and AUC of the significant metrics specific to the numbers task. Significant p-values and AUC over 0.75 are bolded.
AD vs CTL PD vs CTL PD vs PDM
Metrics p-val AUC p-val AUC p-val AUC
@@
Table VII shows the Numbers-specific metrics’ results, studying mostly the variations in time and dimensions between successive answers in the Numbers Task. Of the few metrics that show a significant difference between groups of participants, the time taken to answer is the first metric differentiating between AD and CTL participants, AD participants being slower. In a similar fashion, AD participants are significantly slower over time as compared to CTL participants, an effect of the increased fatigue over the various answers. Finally, we observe that both AD and PD participants tend to input more and more pressure over time compared to CTL participants, once again a potential fatigue indicator.

### v3

@@ III-D1 Grouped metrics
III-E Classification Experiments
In addition to conducting statistical analyses, we conduct binary classification experiments across various experimental groups using metrics extracted from each task. Namely, it involves a binary classification (e.g., AD or CTL) that facilitates the assignment of samples to their corresponding groups, thereby enabling the evaluation of the metrics' effectiveness in distinguishing between the categories under investigation. As for the statistical analysis, the feature set used in the classification models consists of both task-specific and general metrics, ensuring a comprehensive representation of handwriting dynamics. The number of metrics varies across tasks, with Copycube, Drawclock, Copymage Memory, and Copymage containing 54 metrics, while Copyreadtext, Copytext, and Freewrite include 62 metrics. The Numbers task has 68 metrics, whereas Point DOM and Point NONDOM each contain 28 metrics, and Point Sustained has 27 metrics. Additionally, Spiral DOM includes 38 metrics, while Spiral NONDOM and Spiral Pataka each have 37 metrics. These metrics capture various aspects of handwriting, including kinematics (e.g., velocity, acceleration, jerk), spatial properties (e.g., trajectory deviations, area), and pressure dynamics (e.g., mean pressure, pressure variability).
These experiments further evaluate the discriminative power of the metrics and the utility of tasks in extracting valuable information for characterizing different disorders. We are employing a range of classifiers, including Random Forest (RF), Bagging (BG), and Multi-Layer Perceptron (MLP)  [^4] To implement the different classifiers, we are using the Scikit-learn (Sklearn) library in Python. See https://www.tutorialspoint.com/scikit_learn/scikit_learn_introduction.htm. For each classifier, hyperparameter tuning was performed using nested cross-validation (NCV) to ensure optimal generalization performance. Specifically, for the BG classifier, we tuned max_samples (proportion of training samples used per estimator) and n_estimators (number of base estimators). For the MLP classifier, we optimized hidden_layer_sizes (number of neurons in the hidden layer). Finally, for the RF classifier, we tuned n_estimators (number of trees in the forest) and max_metrics (number of metrics considered per split). NCV is a widely used approach for assessing model performance, especially when dealing with limited data, as it helps provide an unbiased estimate of how well the model will generalize to new data. By incorporating two levels of data partitioning, NCV enables both model evaluation and hyperparameter tuning while minimizing overfitting and overly optimistic performance estimates [50, 51]. The NCV framework consists of two levels of cross-validation to ensure robust model evaluation and mitigate overfitting. In our implementation, the outer loop employs a 10-fold cross-validation, partitioning the dataset into 10 folds, with one fold held out for testing while the remaining nine are used for training. Within each outer training set, an inner 10-fold cross-validation is conducted to optimize hyperparameters. The best hyperparameter configuration is selected based on the average F1-score computed across all 10 inner folds, which helps mitigate variance in performance estimates and ensures stability in hyperparameter selection. The F1-score was chosen as the optimization criterion due to class imbalance. To prevent data leakage, cross-validation was strictly performed at the subject level. Speaker-independent folds were created to ensure that all samples from a given individual appeared only in either the training or test set within any iteration. This setup prevents data from the same subject from being in both training and testing, thereby avoiding overly optimistic performance estimates. The entire NCV process involved 100 hyperparameter tuning runs (10 outer folds $\times$ 10 inner folds), with final performance estimates derived by averaging results across the outer folds. This methodology provides a robust assessment of model generalization while maintaining strict speaker independence.
@@ IV-A1 AD vs CTL
Over the 54 task-agnostic metrics, 50 are significant ($p\leq 0.05$) in comparing AD and CTL participants for at least one task, and 29 have p-values smaller than 0.001. We present the 10 metrics having the lowest p-values in Table III across all tasks. Our most significant and consistent metrics across all tasks are the Shannon entropy, the duration, the mean of the speed and accelerations for point tasks, and the variations of angular speed and acceleration for the drawing and writing tasks. We show the box plots for some of those metrics in Figure 9.
Shannon entropy is a measure of the uncertainty in a signal, showing the erratic aspect of one's handwriting, which is a marker of indecision for cognitively impaired participants. The differences measured in linear speed and linear acceleration, as well as duration, all show how significantly slower AD participants are from the control group, presumably due to processing inefficiency. Finally, the differences in angular speed and angular acceleration are indicators of the swiftness of the writing. For detecting AD, the most significant metrics were all for writing and drawing tasks, which are much more demanding cognitively than the rest of tasks and mainly consisting of tasks designed for AD assessment, as opposed to the point and spiral tasks.
@@
A limitation of our approach here is the lower number of participant for the AD/MCI category, which did not allowed us a separated analysis of the cognitive impairments against the participants presenting only AD's signs.
@@
in-air avg. angular acc. 0.070 0.700 <<0.001 0.791 0.087 0.672
in-air shannon entropy <<0.001 0.873 <<0.001 0.887 <<0.001 0.827
in-air std angular acc. 0.019 0.737 <<0.001 0.818 0.034 0.702
in-air std angular speed 0.019 0.730 <<0.001 0.826 0.058 0.692
on-tablet avg. X acc. 0.017 0.732 1.000 0.605 <<0.001 0.761
on-tablet avg. Y acc. 0.078 0.680 1.000 0.597 <<0.001 0.754
@@
in-air avg. angular acc. 0.125 0.691 <<0.001 0.761 <<0.001 0.760
in-air shannon entropy <<0.001 0.794 <<0.001 0.866 <<0.001 0.769
in-air std angular acc. 0.045 0.727 <<0.001 0.768 <<0.001 0.821
in-air std angular speed 0.125 0.694 <<0.001 0.799 <<0.001 0.774
on-tablet avg. X acc. 1.000 0.535 <<0.001 0.788 0.023 0.698
on-tablet avg. Y acc. 1.000 0.597 <<0.001 0.780 0.023 0.700
AD vs CTL Point Dominant Point Non Dominant Point Sustained
@@
in-air avg. angular acc. 0.016 0.726 1.000 0.627 0.030 0.740
in-air avg. acc. <<0.001 0.792 <<0.001 0.850 <<0.001 0.809
in-air avg. speed <<0.001 0.797 <<0.001 0.857 <<0.001 0.827
in-air shannon entropy <<0.001 0.865 <<0.001 0.810 <<0.001 0.872
in-air std angular acc. <<0.001 0.804 0.117 0.708 0.003 0.789
in-air std acc. 0.015 0.716 <<0.001 0.766 0.290 0.655
in-air std speed 0.159 0.650 0.004 0.737 0.390 0.647
@@ IV-A2 PD vs CTL
Over the 54 task-agnostic metrics, 10 are significant ($p\leq 0.05$) in the comparison between PD and CTL participants for at least one task, and 7 have p-values under 0.001. Table IV shows the top 10 metrics with the lowest p-values across all tasks for the PD vs. CTL comparison.
@@
PD vs CTL All Points Point Dominant Point Non Dominant
@@
in-air avg. angular acc. <<0.001 0.723 0.041 0.723 0.009 0.768
in-air avg. acc. 0.027 0.611 1.000 0.611 0.443 0.624
in-air avg. X acc. <<0.001 0.637 0.659 0.628 0.032 0.694
in-air avg. Y acc. <<0.001 0.650 0.076 0.676 0.322 0.647
in-air avg. angular speed <<0.001 0.742 0.021 0.743 0.008 0.768
in-air avg. speed 0.031 0.608 1.000 0.620 0.437 0.630
in-air avg. X speed <<0.001 0.639 1.000 0.616 0.129 0.678
in-air avg. Y speed <<0.001 0.682 0.041 0.711 0.143 0.678
in-air std angular acc. <<0.001 0.723 0.052 0.738 0.010 0.770
in-air std angular speed <<0.001 0.733 0.021 0.756 0.009 0.759
These metrics primarily revolve around in-air dynamics, variations, and averages of speeds and angular speeds. They are notably discriminatory for point tasks. Additionally, we present box plots for these significant metrics in Figure 10 for further visualization and analysis.
@@
For detecting PD, the most significant metrics were for drawing and point tasks. The point tasks were better suited to assess motoric metrics of PD, such as tremor and bradykinesia. Accordingly, we observed subjects with PD showed significantly lower angular speed and angular acceleration in air, which could be indicators of hypometria [52]. The drawing tasks were much more complex tasks, mainly designed for cognitive evaluation but that still require basic motor skills to complete. In those tasks, the in-air angular speed is still a discriminative metric, but metrics of variation (as standard deviation or amplitudes) of the speed and acceleration on the tablet were notable. These findings may represent, tremor, dystonia, dyskinesia, myoclonus other other complicating motor symptoms that develop in some persons with PD [10]. However, none of the task-agnostic metrics measured on the spirals tasks, mainly motor tasks often used for PD assessment, were discriminative enough.
@@ IV-A3 PD vs PDM
Among the 54 task-agnostic metrics, only 6 are significant ($p\leq 0.05$) in comparing PD and PDM participants for at least one task. We only found significant metrics for the Spiral tasks here.
To distinguish PD from other parkinsonism, significant variations can only be found in the Spiral tasks, motor exercises designed for PD assessment. We observe few significant metrics for the Spiral tasks, mostly amplitudes of speed and acceleration, the best being for the vertical velocity. The Spiral tasks reveal a difference in speed between PD and PDM participants, the latest being slower, as shown in Figure 11.
@@ IV-B1 Point Tasks Results
Figure 12 shows all metrics with their significance. None of the metrics are consistently significant across all tasks, but the radius average and radius standard deviation metrics exhibit the same behavior across all tasks: metrics from AD participants are higher than in the other groups. We observe a slightly higher radius average for PD participants compared to CTL, only significant when considering all point tasks, showing that the Point tasks capture only little of the expected tremors.
@@ IV-B2 Spiral Tasks Results
TABLE V: $p$-values and AUC of the specific metrics for all spiral tasks significant for at least one comparison. Significant p-values and AUC over 0.75 are bolded.
Spirals Metric AD vs CTL CTL vs PD PD vs PDM
@@
Spirals N loops 0.092 0.573 <<0.001 0.604 0.303 0.538
avg radius diff 0.141 0.557 0.019 0.568 0.945 0.554
Duration <<0.001 0.695 0.104 0.543 0.003 0.623
Number of Loops 0.370 0.563 <<0.001 0.615 0.209 0.542
Fundamental Period <<0.001 0.694 0.109 0.545 0.004 0.615
Dominant Duration <<0.001 0.695 0.496 0.543 0.155 0.623
Fundamental Period <<0.001 0.694 0.474 0.545 0.185 0.615
Non Dominant N loops 0.714 0.573 0.018 0.604 0.166 0.538
Duration 0.123 0.695 0.043 0.543 0.028 0.623
Number of Loops 0.674 0.563 0.031 0.615 0.241 0.542
Fundamental Period 0.100 0.694 0.051 0.545 0.035 0.615
Pataka N loops 0.150 0.573 0.012 0.604 0.429 0.538
Duration <<0.001 0.695 0.700 0.543 0.105 0.623
Number of Loops 0.406 0.563 0.007 0.615 0.620 0.542
Fundamental Period <<0.001 0.694 0.658 0.545 0.103 0.615
@@ IV-B3 Numbers Task Results
TABLE VI: $p$-values and AUC of the significant metrics specific to the numbers task. Significant p-values and AUC over 0.75 are bolded.
Metric AD vs CTL CTL vs PD PD vs PDM
p-val AUC p-val AUC p-val AUC
@@
Table VI shows the Numbers-specific metrics' results, studying mostly the variations in time and dimensions between successive answers in the Numbers Task. Of the few metrics that show a significant difference between groups of participants, the time taken to answer is the first metric differentiating between AD and CTL participants, AD participants being slower. In a similar fashion, AD participants are significantly slower over time as compared to CTL participants, an effect of the increased fatigue over the various answers. Finally, we observe that both AD and PD participants tend to input more and more pressure over time compared to CTL participants, once again a potential fatigue indicator.